draw the reactant(s)
2 Natty Ph Ph + 2 Na Br + Hz
Answer #1

CH=CH + 2 NaH nate=&nat + H2 12 ph Por Pk-Ç=c-ph + 2Nabr CH=CHT Z NaH +2Ph Br . oh C=c-ph +H₃t ZNaBr. Scanned with CamScanner